OpenStudy (anonymous):

For which value of x does the graph of y = 2x2 − 7x + 6 cross the x-axis? -2/3 -3/2 2 3

hartnn (hartnn):

for the graph to cross the x-axis,its y co-ordinate must be 0 so just put y=0 in the graph equation u will get a quadratic in x solve the quadratic to get x values..

OpenStudy (anonymous):

at x axis,y=0 so, put y=0 0=2x2-7x+6 solving quadratic 2x2-4x-3x+6 2x(x-2)-3(x-2)=0 (2x-3)(x-2)=0 x=3/2,2

OpenStudy (anonymous):

ans matching options=2